The energy of a system increases at a rate of 3.5 t + 6.2 t ², in joules. what is the instantaneous power at t= 3.1 s?


1
Expert's answer
2021-03-24T19:42:22-0400

"W=3.5t+6.2t^2"


"P=dW\/dt=3.5+13t"


"P(3.1)=3.5+13\cdot3.1=43.8\ (W)" . Answer